At 2:00 pm, two cars start toward each other from towns 240 miles apart. If the rate of one car is 10 mph faster than the other, how fast does each car go, if they meet at 5:00 P.M.?
Answer by checkley71(8403)

240/3=x+x+10
2x+10=80
2x=80-10
2x=70
x=70/2
x=35 answer for the slower car.
35+10=45 answer for the faster car.
proof:
3*35+3*45=240
105+135=240
240=240